what is the value of g on the surface of a planet whose mass as well as radius is half of the value for earth
1
Expert's answer
2017-02-23T05:19:25-0500
by definition
g=G*M/R^2 ( G- gravitation constant, M- mass of the planet, R- radius of the planet)

since M1=M/2
and R1=R/2

the g1=g*2

so, new "g" will be 2 times LARGE then original one
if the original g=9.81m/c^2,
the new one g1=19.62m/c^2